OpenStudy (anonymous):

Solve the following system of linear equations by substitution: 6x+3y=30 8x+4y=12

OpenStudy (agreene):

No solutions exist, these are parallel lines.

OpenStudy (anonymous):

Yeah I thought that too

OpenStudy (anonymous):

How can you tell that they are parallel lines?

OpenStudy (agreene):

6x+3y=30 ==> y=10-2x 8x+4y=12 ==> y= 3-2x same slope, different intercepts: parallel, NSE QED.